# Read to Play ([Uncaptioned image] R2-Play): Decision Transformer with Multimodal Game Instruction

### 06 Feb 2024 v1

@@ Abstract
Developing a generalist agent is a longstanding objective in artificial intelligence. Previous efforts utilizing extensive offline datasets from various tasks demonstrate remarkable performance in multitasking scenarios within Reinforcement Learning.However, these works encounter challenges in extending their capabilities to new tasks.Recent approaches integrate textual guidance or visual trajectory into decision networks to provide task-specific contextual cues, representing a promising direction.However, it is observed that relying solely on textual guidance or visual trajectory is insufficient for accurately conveying the contextual information of tasks.This paper explores enhanced forms of task guidance for agents, enabling them to comprehend gameplay instructions, thereby facilitating a “read-to-play” capability.Drawing inspiration from the success of multimodal instruction tuning in visual tasks, we treat the visual-based RL task as a long-horizon vision task and construct a set of multimodal game instructions to incorporate instruction tuning into a decision transformer.Experimental results demonstrate that incorporating multimodal game instructions significantly enhances the decision transformer’s multitasking and generalization capabilities. Our code and data are available at https://github.com/ygjin11/R2-Play
@@ 1 Introduction
To augment the multitasking and generalization capabilities of agents via multimodal game instructions, leveraging a large, diverse offline dataset for pretraining is vital. This approach is supported by the success of instruction tuning in large language models (Ouyang et al.,, 2022; Brown et al.,, 2020) and large multimodal models (Li et al.,, 2023; Liu et al.,, 2023). The Decision Transformer (DT), an innovative integration of transformers with reinforcement learning (RL), demonstrates superior multitasking capabilities in RL, attributed to its large-scale pretraining  (Chen et al.,, 2021; Lee et al.,, 2022). This advancement establishes DT as a leading approach in the domain of large-scale pretraining for RL. Expanding upon the DT framework, we introduce the Decision Transformer with Game Instruction (DTGI). DTGI is pretrained using Multimodal Game Instructions coupled with a comprehensive offline dataset, designed to simulate the agent’s ability to follow game instructions. Simultaneously, we introduce a novel design called SHyperGenerator, which facilitates knowledge sharing between training and unseen game tasks. Experimental results show that integrating multimodal game instruction into DT leads to a substantial improvement in multitasking and generalization performance. Furthermore, multimodal instruction outperforms textual language and visual trajectory, demonstrating its superior capacity to provide detailed and comprehensive context.
@@ 4.2 Baselines
We compare our proposed DTGI with four baselines to answer the questions above.

### 08 Feb 2024 v2

@@ Abstract
Developing a generalist agent is a longstanding objective in artificial intelligence. Previous efforts utilizing extensive offline datasets from various tasks demonstrate remarkable performance in multitasking scenarios within Reinforcement Learning. However, these works encounter challenges in extending their capabilities to new tasks. Recent approaches integrate textual guidance or visual trajectory into decision networks to provide task-specific contextual cues, representing a promising direction. However, it is observed that relying solely on textual guidance or visual trajectory is insufficient for accurately conveying the contextual information of tasks. This paper explores enhanced forms of task guidance for agents, enabling them to comprehend gameplay instructions, thereby facilitating a “read-to-play” capability. Drawing inspiration from the success of multimodal instruction tuning in visual tasks, we treat the visual-based RL task as a long-horizon vision task and construct a set of multimodal game instructions to incorporate instruction tuning into a decision transformer. Experimental results demonstrate that incorporating multimodal game instructions significantly enhances the decision transformer’s multitasking and generalization capabilities. Our code and data are available at https://github.com/ygjin11/R2-Play.
@@ 1 Introduction
To augment agents’ multitasking and generalization capabilities via multimodal game instructions, leveraging a large, diverse offline dataset for pretraining is vital. This approach is supported by the success of instruction tuning in large language models (Ouyang et al.,, 2022; Brown et al.,, 2020) and large multimodal models (Li et al.,, 2023; Liu et al.,, 2023). The Decision Transformer (DT), an innovative integration of transformers with reinforcement learning (RL), demonstrates superior multitasking capabilities in RL, attributed to its large-scale pretraining  (Chen et al.,, 2021; Lee et al.,, 2022). This advancement establishes DT as a leading approach in the domain of large-scale pretraining for RL. Expanding upon the DT framework, we introduce the Decision Transformer with Game Instruction (DTGI). DTGI is pretrained using Multimodal Game Instructions and a comprehensive offline dataset, designed to simulate the agent’s ability to follow game instructions. Simultaneously, we introduce a novel design called SHyperGenerator, which facilitates knowledge sharing between training and unseen game tasks. Experimental results show that integrating multimodal game instruction into DT leads to a substantial improvement in multitasking and generalization performance. Furthermore, multimodal instruction outperforms textual language and visual trajectory, demonstrating its superior capacity to provide detailed and comprehensive context.
@@ 4.2 Baselines
We compare our proposed DTGI with four baselines.